Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (howard-wolowitz):

question in comments, blood type question

OpenStudy (4everaddicted2anime):

if you add all of the probabilities it should equal 1.00 Add the one you were given and subtract that from 1.00

OpenStudy (4everaddicted2anime):

*ones

OpenStudy (howard-wolowitz):

ok so the answer is 0.25

OpenStudy (universetoast):

.5+.2+.05=.75 1.00-.75=.25 The decimals are based on percentages 1.00=%100

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!